Simple Function Evaluation:
f(x)=1/3x5(power of 5), find f(3)
Please help. I cannot get this one!

Answer by tutorcecilia(2152)

f(x)=1/3x^5
f(3)=1/3(3)^5 [substitute (3) for (x)]
f(3)=1/(3)(243) [Simplify]
f(3)=1/729